# Revealing Multimodal Contrastive Representation Learning through Latent Partial Causal Models

### v1

@@ A.5 The Proof of identifiability on convex bodies
$-2\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\big{[}e^{d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}.$ (50)
We then connect the right-hand side of Eq. (41) with Eq. (50). To this end, since the two terms in the right-hand side of Eq. (41) are symmetrical, we focus on one of the two terms for convenience, e.g., $\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$. It can be shown that:
$\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$ (51)
$\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}\operatorname*{\mathbb{E}}\limits_{\tiny\mathbf{z}_{t}\sim p(\mathbf{z}_{t}|\mathbf{z}_{x})}[-\log q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})]\big{]}}$ (52)
$=$ $-\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-d(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau+\log C_{q}(\mathbf{z}_{x})\big{]}}$ (53)
$=$ $\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-d(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau\big{]}+\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x})\sim p(\mathbf{z}_{x})}[\log C_{q}(\mathbf{z}_{x})]}$ (54)
$=$ $\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}-d(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau\big{]}+\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x})\sim p(\mathbf{z}_{x})}[\log\int e^{(d(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau)}\mathrm{d}\mathbf{z}_{x}]}$ (55)
@@
$=-\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}d(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}$ (56)
@@
$\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{t})\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t})),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}=-\operatorname*{\mathbb{E}}\limits_{\tiny(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\big{[}e^{d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}.$ (57)
@@
A.5.2 Identifiability result on Convex Bodies
Theorem 3.2 represents a symmetrical adaptation of Theorem 3 from Zimmermann et al. (2021). This alignment allows us to employ Propositions 4, Lemma 1 and Lemma A from Zimmermann et al. (2021) to demonstrate that the global minimization of the objective outlined in Equation (41), as specified in Theorem 3.2, identifies the latent variables $\mathbf{z}_{x}$, as well as $\mathbf{z}_{x}$, up to linear transformations. For completeness, a brief proof is provided herein, with comprehensive details available in the original work. Clearly, the global minimum of the cross-entropy between two distributions is reached if they match by value and have the same support. Therefore, for the optimal solution of the objective loss Eq. (14) in Theorem 3.2, we have:
@@ A.5.2 Identifiability result on Convex Bodies
$C_{p}(\mathbf{z}_{x})e^{-\delta(\mathbf{z}_{x},\mathbf{z}_{x})/\lambda}=C_{q}(\mathbf{z}_{x})e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{x}))/\tau},$
@@
Then, by limiting $\delta$ be an $L^{\alpha}$ metric for $\alpha\geq 1$, $\alpha\neq 2$ or the $\alpha$-th power of such an $L^{\alpha}$ metric, using the Theorem 5 in Zimmermann et al. (2021), Theorem 3.1 can identify the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ up to an permutation transformation, i.e., the recovered latent variable $\mathbf{\hat{z}}_{x}$, obtained through the minimization of Equation (11), is related to the true $\mathbf{z}_{x}$ as follows: $\mathbf{\hat{z}}_{x}=\mathbf{P}\mathbf{z}_{x}+\mathbf{c}$, where $\mathbf{P}$ represents an permutation matrix with scaling, and $\mathbf{c}$ is a constant vector.
A.6 More Results on CelebA
(a) Smile
(b) Brightness
(c) Hair color
(d) Eye shadow
(e) Gender w/ Mustache
(f) Glasses
Figure 3: Disentangled Representations learned by combining pre-train CLIP and FastICA.
(a) Face width
(b) Face Size
(c) Background color
(d) Skin color
(e) Azimuth
(f) Background remove
Figure 4: Disentangled Representations learned by combining pre-train CLIP and FastICA.
(a) Face Height
(b) Hair Style
(c) Hair length
(d) Age
Figure 5: Disentangled Representations learned by combining pre-train CLIP and FastICA.
A.7 More Results on ImageNet-Type Data
@@
A.8 Implementation Details
To obtain disentangled representations for the CelebA dataset, we initially employ the FastICA implementation available in the scikit-learn software on the features extracted from the pretrained ViT-B/32 encoder. Subsequently, we train the decoder, as outlined in Table 8, utilizing Mean Squared Error (MSE) loss.
@@ A.8 Implementation Details
In our experiments with ImageNet-Type data, we utilized the PCA and FastICA implementations provided by scikit-learn. For our proposed method, which combines PCA and ICA, we configured the number of components to 500 for PCA, and for FastICA, we set it to 160 for 1, 2, and 4-shot learning scenarios, and 200 for 8 and 16-shot learning scenarios. When employing ICA alone, we chose to use 300 components. For the proposed method with ICA only, we set number of components to 300.
Following the setting of linear probe in CLIP, we train a logistic regression classifier using scikit-learn’s L-BFGS implementation, with maximum 1,000 iterations. We determine the L2 regularization strength using a hyperparameter sweep on the validation sets over the range between $10^{-6}$ and $10^{6}$ , with 96 logarithmically spaced steps. To save compute required for the sweeps, we perform a parametric binary search and iteratively halves the interval around the peak until it reaches a resolution of 8 steps per decade. The hyperparameter sweeps are performed on a validation split of each dataset.

### v2

@@ Appendix E The Proof of identifiability on convex bodies
$2\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}$ $+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{-d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}$
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\big{[}e^{-d\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}.$ (37)
We then connect the right-hand side of Eq. (32) with Eq. (37). To this end, since the two terms in the right-hand side of Eq. (32) are symmetrical, we focus on one of the two terms for convenience, e.g., $\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$. It can be shown that:
$\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}H(p(\mathbf{z}_{t}|\mathbf{z}_{x})),q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x}))\big{]}}$ (38)
$=$ $\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}{\big{[}\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t}|\mathbf{z}_{x})}[-\log q_{\mathbf{h}}(\mathbf{z}_{t}|\mathbf{z}_{x})]\big{]}}$ (39)
$=$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau+\log C_{q}(\mathbf{z}_{x})\big{]}}$ (40)
$=$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau\big{]}+\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x})\sim p(\mathbf{z}_{x})}[\log C_{q}(\mathbf{z}_{x})]}$ (41)
$=$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau\big{]}+\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x})\sim p(\mathbf{z}_{x})}[\log\int e^{(-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau)}\mathrm{d}\mathbf{z}_{x}]}$ (42)
@@
$=\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t}))/\tau\big{]}}+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\big{[}e^{-\delta\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}+\log|\mathcal{Z}_{c}|$ (43)
@@
$\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{t})\sim p(\mathbf{z}_{t})}{\big{[}H(p(\mathbf{z}_{x}|\mathbf{z}_{t})),q_{\mathbf{h}}(\mathbf{z}_{x}|\mathbf{z}_{t}))\big{]}}=$ $\operatorname*{\mathbb{E}}\limits_{(\mathbf{z}_{x},\mathbf{z}_{t})\sim p(\mathbf{z}_{x},\mathbf{z}_{t})}{\big{[}\delta\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau\big{]}}$ (44)
$+\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{t}\sim p(\mathbf{z}_{t})}\Big{[}\log\operatorname*{\mathbb{E}}\limits_{\mathbf{z}_{x}\sim p(\mathbf{z}_{x})}\big{[}e^{-\delta\big{(}\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{t})\big{)}/\tau}\big{]}\Big{]}+\log|\mathcal{Z}_{c}|.$ (45)
@@
E.2 Identifiability result on Convex Bodies
Theorem 4.3 represents a symmetrical adaptation of Theorem 3 from [67]. This alignment allows us to employ Propositions 4, Lemma 1 and Lemma A from [67] to demonstrate that the global minimization of the objective outlined in Eq. (32), as specified in Theorem 4.3, identifies the latent variables $\mathbf{z}_{x}$, as well as $\mathbf{z}_{x}$, up to linear transformations. For completeness, a brief proof is provided herein, with comprehensive details available in the original work. Clearly, the global minimum of the cross-entropy between two distributions is reached if they match by value and have the same support. Therefore, for the optimal solution of the objective loss Eq. (10) in Theorem 4.3, we have:
@@ E.2 Identifiability result on Convex Bodies
$C_{p}(\mathbf{z}_{x})^{-1}e^{-\delta(\mathbf{z}_{x},\mathbf{z}_{x})/\lambda}=C_{q}(\mathbf{z}_{x})^{-1}e^{-\delta(\mathbf{h}(\mathbf{z}_{x}),\mathbf{h}(\mathbf{z}_{x}))/\tau},$
@@
Then, by limiting $\delta$ be an $L^{\alpha}$ metric for $\alpha\geq 1$, $\alpha\neq 2$ or the $\alpha$-th power of such an $L^{\alpha}$ metric, using the Theorem 5 in [67], Theorem 4.3 can identify the latent variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ up to an permutation transformation, i.e., the recovered latent variable $\mathbf{f}_{x}(\mathbf{x})$, obtained through the minimization of Eq. (8), is related to the true $\mathbf{z}_{x}$ as follows: $\mathbf{f}_{x}(\mathbf{x})=\mathbf{P}\mathbf{z}_{x}+\mathbf{c}$, where $\mathbf{P}$ represents an permutation matrix with scaling, and $\mathbf{c}$ is a constant vector.
Appendix F Differences from Previous Analysis for Mutilmodal Contrastive Learning
This work differs from previous works focusing on identifiability analysis for multimodal settings [8, 62, 14] across the following three key dimensions.
Modeling Setting
This work proposes modeling transferable knowledge across modalities by latent coupled variables. In contrast, previous works [8, 62, 14] often achieve this by introducing the same/shared variables. The advantages of employing latent coupled variables are thoroughly justified in Section 2. Loosely speaking, From the perspective of model flexibility, the proposed model can be considered a generalization of previous works. This generalization is apparent as the proposed model seamlessly reduces to a single-modal setting when the mixing functions from latent space to observed space are enforced to be identical, and specific variables are omitted.
Identifiability Results
The identifiability results obtained in this work diverge from those found in previous works [8, 62], both in terms of breadth and depth of identifiability, due to the introduction of the undirected edge between $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$. a) Breadth of Identifiability: Unlike earlier works that often achieve only partial identifiability of latent coupled variables $\mathbf{z}_{x}$ or $\mathbf{z}_{t}$, e.g., latent content variables but not latent style variables [8, 62], our model extends this scope to ensure complete identifiability of latent coupled variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$. b) Depth of Identifiability: In terms of depth, this work identifies latent coupled variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$ up to linear or permutation transformations. As a result, after applying a linear ICA method, we can obtain component-wise identifiability, i.e., recovering independent latent variables up to permutation and scaling. This level of precision offers an enhancement over the block identifiability result in previous studies [8, 62], which only identifying latent variables up to a nonlinear invertible mapping, even for independent latent variables. The differences above in both breadth and depth of identifiability results enable us, for the first time, to unveil the component-wise disentanglement capabilities of multimodal contrastive representation learning.
Practical Significance in Real Applications
Prior studies [8, 62, 14] have predominantly relied on simulation experiments, which often encounter a substantial gap between the assumptions made in theoretical analyses and the practical conditions of real-world applications. In contrast, our work bridges this gap by validating our theoretical findings using pre-trained CLIP models on over 16 diverse real-world datasets. This empirical approach not only substantiates the practical effectiveness of our theoretical results but also demonstrates their applicability and robustness in real-world multimodal settings, highlighting a significant departure from previous work in terms of real-world applicability.
Appendix G Differences from Previous Analyses for Single-Modal Contrastive Learning
This work sets itself apart from prior studies focused on the analysis of single-modal contrastive learning [67, 58] in the following key aspects.
Problem Context
Previous works primarily address single-modal scenarios, whereas our proposed model extends this framework to the more complex multimodal domain. This extension can be viewed as a generalization of prior approaches. Specifically, our model naturally reduces to a single-modal setting when the mixing functions from the latent space to the observed space are identical, and certain variables are omitted. By expanding the scope to multimodal data, our approach addresses the limitations of prior studies and provides a more comprehensive understanding of contrastive learning.
Technical Perspective
Addressing multimodal settings requires significantly broader technical developments compared to single-modal analyses. To this end, we developed Theorem 3.1, which generalizes the asymptotic analysis of contrastive learning to the multimodal context, providing a robust theoretical foundation. Bridging the gap between single-modal and multimodal settings also necessitated novel theoretical insights. For instance, Theorem 4.1 and Theorem 4.3 establish critical connections between multimodal contrastive learning and traditional single-modal frameworks, enabling a unified understanding across these domains. These results not only expand the applicability of contrastive learning but also highlight the intricate dependencies introduced by multimodal data.
New Insights
In the multimodal context, a key challenge is effectively modeling the connections between different modalities. This motivates the central insight of our work: latent coupled variables, linked by a unidirectional edge, provide a foundation for exploring whether partial causal models are sufficient for multimodal learning. As highlighted in the introduction, we offer theoretical support for the success of multimodal contrastive learning, including guarantees for its disentanglement capabilities. From a practical perspective, we recommend refining representations from pre-trained CLIP-like models rather than using them directly. Specifically, applying linear ICA methods, such as FastICA (aligned with assumptions on the hypersphere), or combining PCA and FastICA (aligned with assumptions on convex bodies), can enhance performance on tasks that rely on disentangled representations. These insights not only validate the robustness of our theoretical findings but also emphasize their practical significance in real-world applications.
Appendix H More Results on CelebA
Figures LABEL:fig:_celeba11 - LABEL:fig:_celeba1 illustrate the 16 distinct disentangled representations obtained using pre-trained CLIP with FastICA. Interestingly, our method achieves competitive results compared to specialized disentanglement techniques, such as FactorVAE [33] and $\beta$-TCVAE [6]. Specifically, FactorVAE identified 8 disentangled attributes, while $\beta$-TCVAE reported 15, whereas our approach successfully discerns 16 distinct disentangled representations.
It is important to note that this comparison is not meant to position our method as a more effective disentanglement technique. Rather, our experiments are designed solely to validate our theoretical findings. We present this comparison to provide insight into the potential of leveraging CLIP for learning disentangled representations, thereby motivating future research in this direction. A particularly interesting avenue could be exploring how disentanglement capabilities relate to the manipulation of pre-trained vision models, such as diffusion models.
Appendix I More Results on ImageNet-Type Data
@@
Source Target (ImageNet-)
Encoders Methods ImageNet V2 Sketch R A Avg.
RN50 Lin. P. 41.34 33.67 11.55 26.27 9.67 20.29
Lin. P. w/ FastICA 44.10 36.07 12.75 30.15 11.64 22.65
Lin. P. w/ PCA and FastICA 42.86 35.38 12.29 28.81 9.79 21.57
RN101 Lin. P. 48.23 39.53 18.80 38.10 14.32 27.69
Lin. P. w/ FastICA 49.43 41.02 17.49 39.33 15.25 28.27
Lin. P. w/ PCA and FastICA 49.01 40.25 19.26 39.71 14.75 28.49
ViT32 Lin. P. 47.82 39.53 21.51 40.94 15.99 29.49
Lin. P. w/ FastICA 49.43 40.66 22.66 41.78 16.41 30.38
Lin. P. w/ PCA and FastICA 49.48 41.09 23.72 43.48 16.77 31.27
ViT16 Lin. P. 54.30 46.06 27.58 50.76 29.24 38.41
Lin. P. w/ FastICA 56.65 48.18 28.27 55.50 33.39 41.33
Lin. P. w/ PCA and FastICA 56.16 47.46 30.21 55.49 31.71 41.22
@@
Appendix J More Results on Few-Shot Learning Task
Appendix K Implementation Details
We perform all experiments using the GPU RTX 4090, equipped with 32 GB of memory.
Synthetic Data
We consider latent coupled variables $\mathbf{z}_{x}$ and $\mathbf{z}_{t}$, each with a dimensionality of 10. Additionally, we have modality-specific latent variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$, both set to a dimension of 5. The process begins with sampling from the marginal distribution $p(\mathbf{z}_{x})$, and the samples of modality-specific latent variables $\mathbf{m}_{x}$ and $\mathbf{m}_{t}$ are obtained by sampling from Gaussian distributions with zero mean and one variance. We then create real pairs by sampling from the conditional distribution $p(\mathbf{z}_{t}|\mathbf{z}_{x})$. The observational data $\mathbf{x}$ and $\mathbf{t}$ are generated using two different Multi-Layer Perceptrons (MLPs). Specifically, we utilize MLPs comprising three hidden layers with leaky ReLU units and random weights. To ensure the invertibility of the MLP g, we carefully control the condition number of the weight matrices. For our encoders concerning both $\mathbf{z}_{t}$ and $\mathbf{z}_{x}$, we adopt an MLP architecture with leaky ReLU units.
Evaluation:
To evaluate the linear identifiability result established in Corollary 4.2, we assess how well the learned representations $\mathbf{\hat{z}}_{x}$ preserve the structure of the ground-truth latent variables $\mathbf{z}_{x}$ up to a linear transformation. Specifically, we perform the following steps:
  1. 1.
    Learned Representations Extraction: We first obtain representations $\mathbf{\hat{z}}_{x}$ learned by multimodal contrastive learning.
  2. 2.
    Linear Regression Fitting: We fit a linear regression model of the form:
    $\mathbf{\hat{z}}_{x}=\mathbf{\hat{A}}\mathbf{z}_{x}+\mathbf{\hat{c}}+\epsilon,$
    where $\mathbf{\hat{A}}$ is a learned transformation matrix, $\mathbf{\hat{c}}$ is an offset vector, and $\epsilon$ represents residual errors.
  3. 3.
    Coefficient of Determination ($R^{2}$) Computation: We compute the $R^{2}$ score, defined as:
    $R^{2}=1-\frac{\sum_{i}\|\mathbf{\hat{z}}_{x,i}-(\mathbf{\hat{A}}\mathbf{z}_{x,i}+\mathbf{\hat{c}})\|^{2}}{\sum_{i}\|\mathbf{\hat{z}}_{x,i}-\bar{\mathbf{\hat{z}}}_{x}\|^{2}},$
    where $\bar{\mathbf{\hat{z}}}_{x}$ is the mean of $\mathbf{\hat{z}}_{x}$. This metric quantifies how well the learned representations can be linearly mapped to the true latent variables.
  4. 4.
    Analysis Under Different Assumption Violations: We repeat the evaluation under settings that both satisfy and violate the theoretical assumptions, as listed in Table 3, allowing us to empirically assess the robustness of the identifiability results.
By reporting the $R^{2}$ scores across different conditions, we quantify the extent to which multimodal contrastive learning successfully recovers the latent variables up to a linear transformation.
to evaluate permutation identifiability result in Corollary 4.4, we employ the mean correlation coefficient (MCC) between the ground-truth $\mathbf{z}_{x}$ and representations $\mathbf{f}_{x}(\mathbf{x})$ learned by multimodal contrastive learning. To compute MCC, we follow these steps:
  1. 1.
    Compute Correlation Coefficients: We first calculate the correlation coefficients between all pairs of ground-truth source variables and representations learned by multimodal contrastive learning. Specifically, for each pair of source component $\mathbf{z}_{x,i}$ and recovered latent component $\mathbf{\hat{z}}_{x,j}$, we compute the Pearson correlation coefficient:
    $\rho_{i,j}=\frac{\text{Cov}(\mathbf{z}_{x,i},\mathbf{\hat{z}}_{x,j})}{\sigma_{\mathbf{z}_{x,i}}\sigma_{\mathbf{\hat{z}}_{x,j}}},$ (49)
    where $\text{Cov}(\cdot,\cdot)$ denotes the covariance, and $\sigma_{\mathbf{z}_{x,i}}$ and $\sigma_{\mathbf{\hat{z}}_{x,j}}$ are the standard deviations of the respective components.
  2. 2.
    Solve the Linear Sum Assignment Problem: Since the estimated components may be permuted relative to the ground-truth variables, we solve a linear sum assignment problem to determine the optimal one-to-one mapping between the ground-truth and the learned representations. The goal is to maximize the total absolute correlation across all assigned pairs.
  3. 3.
    Compute the Mean Correlation Coefficient (MCC): Given the optimal assignment of the ground-truth variables to the learned representations, we compute the mean of the absolute values of the assigned correlation coefficients:
    $\text{MCC}=\frac{1}{d}\sum_{i=1}^{d}|\rho_{i,\pi(i)}|,$ (50)
    where $\pi(i)$ denotes the index of the assigned representation corresponding to the $i$th latent variable, and $d$ is the number of latent variables.
A high MCC indicates that the learned representation closely match the true source variables, up to permutation transformations, thereby validating the identifiability of the learned representations.
@@
Disentangled Representation Learning on CelebA
To obtain disentangled representations for the CelebA dataset, we initially employ the FastICA implementation available in the scikit-learn software on the representations extracted from the pretrained ViT-B/32 encoder. Subsequently, we train the decoder, as outlined in Table 8, utilizing Mean Squared Error (MSE) loss.
Experiments of Linear Probe
In our experiments with ImageNet-Type data, we utilized the PCA and FastICA implementations provided by scikit-learn. For our proposed method, which combines PCA and ICA, we configured the number of components to 500 for PCA, and for FastICA, we set it to 160 for 1, 2, and 4-shot learning scenarios, and 200 for 8 and 16-shot learning scenarios. When employing ICA alone, we chose to use 300 components. For the proposed method with ICA only, we set number of components to 300. Following the setting of linear probe in CLIP, we train a logistic regression classifier using scikit-learn’s L-BFGS implementation, with maximum 1,000 iterations. We determine the L2 regularization strength using a hyperparameter sweep on the validation sets over the range between $10^{-6}$ and $10^{6}$ , with 96 logarithmically spaced steps. To save compute required for the sweeps, we perform a parametric binary search and iteratively halves the interval around the peak until it reaches a resolution of 8 steps per decade. The hyperparameter sweeps are performed on a validation split of each dataset.
FastICA as a plug-and-play Tool.
We incorporate FastICA in the framework proposed in [63] to enhance its ability for few shot learning. The framework consists of two primary modules: one keeps the zero-shot capabilities of pre-trained CLIP, ensuring effective utilization of prior knowledge, while the other, the cache module, constitutes the central contribution of the work. The cache module endeavors to transfer knowledge from labeled training samples. Given the above, we integrate FastICA into the cache module, preserving the invaluable prior knowledge derived from the zero-shot abilities of pre-trained CLIP. For parameter settings in FastICA, we opted for 100 components for the majority of datasets. Specifically, we assigned 350 components for the ImageNet dataset, 300 components for the OxfordPets dataset, and 50 components for the EuroSAT dataset. A learning rate of 0.1 was employed for implementation. For the remaining parameter settings, we adhered to the specifications outlined by [63].